We're going to leave birds for a minute and work through some probability calculations using a coin.

A coin has two sides-head and tail.

When we talk about probability we use numbers between 0 and 1. An event with a probability of 0 will never happen. An event with a probability of 1 will definitely happen.

If I toss a coin what is the probability it will land heads up?

a. 0

b. 1/4

c. 1/2

d. 1
Answer #1

for a coin there are 2 sides, one is head and one is tail so when we flip a coin there are 2 possibilities, we can get either a head or a tail, that is there is an equal chance of getting head or tail, that is the probability of getting head =1/2=probability of getting tail.

so the probability of getting head= number of head/total number

=1/2

c) 1/2